OpenStudy (anonymous):

f:(0;+∞)->R,f(x)=ln(1+1/x) Calculate the limit of the string:

OpenStudy (anonymous):

\[(An)n \ge1 ;An=f(1)+f(2)+...+f(n)-\ln(2n+1)\]

ganeshie8 (ganeshie8):

telescoping series

ganeshie8 (ganeshie8):

\(\large \sum \limits_{x=1}^n f(x) \) \(\large \sum \limits_{x=1}^n \ln(1 + \frac{1}{x}) \) \(\large \sum \limits_{x=1}^n \ln(\frac{x+1}{x}) \) \(\large \sum \limits_{x=1}^n \ln(x+1) - \ln (x) \) \(\large [\ln(1+1) - \ln (1) ] + [\ln(2+1) - \ln (2) ] + ... [\ln(n+1) - \ln (n) ] \) \(\large [\ln(2) - \ln (1) ] + [\ln(3) - \ln (2) ] + ... [\ln(n+1) - \ln (n) ] \)

ganeshie8 (ganeshie8):

everything in the middle cancels out, leaving u wid : \(\large -\ln(1) +\ln(n+1)\)

OpenStudy (anonymous):

an this row tends to....................?

ganeshie8 (ganeshie8):

since ln(1) = 0, that gives u \(\large \ln(n+1)\)

OpenStudy (anonymous):

yeah i got it thanks

ganeshie8 (ganeshie8):

so, \(\sum \limits_{x=1}^n \ln(1 + \frac{1}{x}) = \ln(n+1) \)

ganeshie8 (ganeshie8):

\(An=f(1)+f(2)+...+f(n)-\ln(2n+1)\) \(An=\ln(n+1)-\ln(2n+1)\) \(An=\ln(\frac{n+1}{2n+1})\)

ganeshie8 (ganeshie8):

u wlc :)